Journalist Report – November 2nd

by Marina Buqueras Rovira


Streaming y Film en Marte

Sol 4

9 am Testeando las comunicaciones con la Tierra, todo funciona bien con la TV BDN!
La Universidad de Purdue llegó y David nuestro Commander los recibió en nuestra casa.

10 am Un buen desayuno de pancakes con arándanos y chocolate nos hace perder la noción.

12 pm La tripulación se pone de gala y ya estamos listos para la entrevista en vivo de la televisión Badalona Comunicació. La buena vibra nos mantiene con humor, el Commander y el XO, hacen que la Crew 285 sea un equipo de diez!

14 pm La sopa caliente nos recompone de la ardua sensación de nostalgia, después de nuestro directo, ya que sabemos que nuestros familiares nos han visto y esperemos que les lleguen nuestros abrazos.

16 pm La tarde, en el SciencieDome, la dedicamos a nuestros proyectos, Andrea sigue con sus mezclas de materiales de Marte para su experimento y Luis preparándose para su próxima observación solar y su astrofotografía del día, Andrés manteniendo y plantando nuevas semillas en el GreenHab, Marina les acompaña para captar todos estos momentos.

Tom, hoy fue el nuevo actor de la MDRS, nuestro Engineer és el nuevo Matt Damon en The Martian!

Mañana nos espera un día intenso con gran actividad, así que acabamos el sol 4 preparando bien los EVA.

Seguimos fuertes!


Streaming and Film on Mars

Sol 4

9 am Testing communications with Earth, everything is working well with TV BDN!
Purdue University has arrived, and our Commander David welcomed them to our home.

10 am A delicious breakfast of blueberry and chocolate pancakes makes us lose track of time.

12 pm The crew dresses up, and we are all set for the live interview on Badalona Comunicació TV. The good vibes keep us in high spirits; the Commander and the XO make Crew 285 a top-notch team!

2 pm Hot soup revitalizes us from the intense feeling of nostalgia after our live session, knowing that our families have watched, and we hope they feel our hugs reaching them.

4 pm The afternoon at the ScienceDome is dedicated to our projects. Andrea continues with her experiments with Martian materials, Luis prepares for his next solar observation and daily astrophotography, Andrés maintains and plants new seeds in the GreenHab, and Marina comes with them to capture all these moments.

Tom, today, took on the role of the new actor at MDRS; our Engineer is the new Matt Damon in "The Martian"!

Tomorrow, we anticipate a busy day with significant activity, so we conclude Sol 4 by preparing well for the EVAs.

We remain strong!

Supplemental Operations Report – November 5th

Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Rovers: Perseverance, Curiosity, Spirit, Sojourner. A tunnel stick on the right from the rear airlock.

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 19 %; Avg 54.4%.
VDC Last 24 hours: Max 58.58V; Min 43.59V; Avg 50.65V.
Generator run time: 3928.8 hours. Oil changed on 11/05/2023

Propane Readings:
Station Tank: 79%
Director Tank: 82%
Intern Tank: 85%
Generator Tank: 80%"

Hab Static Tank – 347 gallons
GreenHab – 168 gallons
Outpost tank – 350 gallons
Science Dome – 0 gallons
Hab Toilet Tank emptied: Yes

Sojourner rover used: No.
Hours: 195.4
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: All rovers still waiting for all parts to arrive.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: To Hanksville for supplies.
General notes and comments: All cars awaiting renewed insurance ID.

Summary of Internet: All nominal.
EVA suits and radios:
Suits: All nominal
Comms: All nominal

Campus wide inspection, if action taken, what and why: One tunnel stick on the right from the rear airlock is corroded out at the ground level and needs to be replaced.
Summary of Hab Operations: Toilet tank level sensors need to be cleaned from toilet paper.
Summary of GreenHab Operations: All nominal
Summary of SciDome Operations: All nominal
Summary of Observatories Operations: Robotic observatory had a mount error but currently is operational. Solar observatory but nominal.
Summary of RAM Operations: Remote temperature sensor for smart home systems is offline. Most likely due to the battery being dead and it needs to be relocated to avoid touching metal walls.
Summary of Outpost Operations: Directors trailer thermostat was randomly offline three times. A new one is ordered to diagnose this issue.
Summary of Health and Safety Issues: All nominal.

Astronomy Report – November 3rd [status draft



Name: Luis Diaz
Crew: 285
Date: 11/03/2023
Robotic Telescope Requested: MDRS-WF
Objects to be Imaged this Evening: Accumulate more light for Helix Nebula, Horsehead Nebula and maybe others.
Images submitted with this report: Pacman Nebula (NGC 281)
Problems Encountered: –
Solar Features Observed: –
Images submitted with this report: –
Problems Encountered: –

GreenHab Report – November 2nd

GreenHab Officer: Andres David Reina
Environmental control: (heater, fan, door open?)
Open door: 9:30 am:
-     Temperature: 23,4C
-     Humidity:10%
Close door: 18:05
-     Temperature: 21,9C
-     Humidity: 11%

Hours of supplemental light: None
Daily water usage for crops: 15 gallons.
Daily water usage for research and/or other purposes: None
Water in Blue Tank 177.94 gallons
Time(s) of watering for crops: n/a
Changes to crops: A slight growth in the tomato crop. Two existing germinations with an average size of 2.5 cm in height. Growth is observed in cucumber and lettuce crops.

– The Greenhab door opened at 9:30 am. Environmental measurements of temperature and relative humidity were taken. The GreenHab was checked and watered. Germination can be evidenced in the cucumber crop.The average temperature and humidity on the day was approximately 23.55 C and 10.5%. The crops began to be watered with 15 gallons according to the instructions received.
Support/supplies needed: n/a

Astronomy Report – November 2nd



Name: Luis Diaz
Crew: 285
Date: 11/02/2023
Robotic Telescope Requested: MDRS-WF
Objects to be Imaged this Evening: Helix Nebula and maybe others.
Images submitted with this report: Pleiades (M45)
Problems Encountered: –
Solar Features Observed: –
Images submitted with this report: –
Problems Encountered: –

Sol Summary – November 1st

Sol: 3
Summary Title: Resting Day
Author’s name: David Mateus
Mission Status: Nominal
Sol Activity Summary: The crew started the day at 9 a.m., after a deserved rest. On Sol 2, the crew worked hard and ended the day with a small “Halloween party” with pizza, mocktails, and Latin music.
The day was focused on the projects, preparing future EVAs
Look Ahead Plan: Continue with the project planning of the EVAs and be prepared for tomorrow´s media events.
Anomalies in work: None
Weather: Partly cloudy
Crew Physical Status: Nominal
EVA: No EVA today
Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, Astronomy report.
Support Requested: We need to add a new email address to the mailing list of mission support because our HSO is having issues sending emails from her current email to mission support; the new email we need to add is andrea10stars1g. For the media stream with Badalona Comunicacio, I informed the Media and Public Relations Director and MDRS Director we need an internet connection tomorrow between 8 a.m. and 2 p.m.

Operations Report – November 1st

SOL: 3
Name of person filing report: Tomás Ignacio Burroni
Non-nominal systems: None
Notes on non-nominal systems: N/A
Spirit rover used: No
Currently charging: Yes
Opportunity rover used: No
Currently charging: Yes
Curiosity rover used: No
Currently charging: Yes
Perseverance rover used: No
Currently charging: Yes
General notes on rovers: N/A
Summary of Hab operations:
WATER USE: 23 gallons
Water (static tank): 406 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): off
Toilet tank emptied: yes, the indicator only got to “⅓” instead of “empty”
Summary of internet: No comments
Summary of suits and radios: No comments
Summary of GreenHab operations: fertilization, thermostat adjusted to 65 F
WATER USE: 3.6 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: used the blender, 3 large and 1 small paper filters, the microwave, the oven, 2 large erlenmeyer flasks, 1 large and 4 small beaker glasses, the scale, and the mortar and pestle. All this was used to recover starch from rice and potatoes and then mix it with a soil sample to create a concrete-like material.
Dual split: Off
Summary of RAM operations: Not used today
Summary of any observatory issues: “Mount error” was fixed by the observatory director, however 2 hs later the “Camera error” showed up again, which stopped the image acquisition.
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: None

Journalist Report – November 1st

by Marina Buqueras Rovira


¡El que persevera, alcanza!

Sol 3

La noche terminó con los sentimientos al máximo, el agotamiento nos arropaba aunque la satisfacción de lo conseguido nos hacía vibrar de felicidad, ¡las risas nos hacen hasta llorar!

¡Amaneció tranquilo…, el amanecer aquí es tan bonito!

9:30 am Desayunamos y empezamos a poner en marcha cada uno de los trabajos planificados para hoy.

La limpieza general en el Hab entre todos, Andrés acabando el inventario de los alimentos (nos tiene las galguerías, golosinas, dulces, chucherías… bien restringidas), Luis concentrado y obteniendo imágenes inimaginables del cielo diariamente, Andrea se prepara para iniciar las mezclas y obtener el almidón en el ScienceDome, Tomás con los receptores gps…

La mañana pasa muy rápido, tan concentrados…

14:30 pm Tom y Marina cocinaron a dúo el almuerzo para la tripulación, hoy tenemos brócoli con boniato y zanahoria acompañado de puré de patatas. Y el momento sobremesa nos recompone el alma.

17 pm La tarde siguió con calma pero sin pausa, Andrés se dirige al Greenhab a fertilizar y los demás enfocados en sus tareas y proyectos.

Mañana más y mejor!


The one who perseveres, achieves!

Sol 3

The night ended with emotions running high, exhaustion enveloping us even as the satisfaction of our achievements made us thrilled with happiness – laughter making us even cry!

The morning dawned peacefully; the sunrise here is so beautiful!
9:30 am We had breakfast and began to set each of today’s tasks in motion.

General cleaning in the Hab, Andrés finishing the food inventory (he has our snacks, treats, sweets, and candies well rationed), Luis focused on capturing unbelievable images of the sky daily, Andrea is getting ready to start the mixtures and obtain starch in the ScienceDome, Tomás with the GPS receivers…

The morning goes by so fast, so focused…

2:30 pm Tom and Marina cooked lunch together for the crew. Today we had broccoli with sweet potato and carrot accompanied by mashed potatoes. And the post-meal moment reassembles our souls.

5 pm The afternoon continued calmly but without pause, Andrés headed to the Greenhab to fertilize, and the others focused on their tasks.

More and better tomorrow!

Copyright © The Mars Society. All rights reserved. | Main Site